Solve the following system of equations.
3x + 2y = -12 and x = 2

Answer by mananth(16949)

3x + 2y = -12 and x = 2
plug the value of 2 in the equation
3x + 2y = -12
3*2 +2y=-12
6+2y=-12
2y=-18
y=-9